OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? y = 4x – 2 y = 2x – 4 y = –3(x – 4) y = 3(x + 4)

OpenStudy (anonymous):

y = mx + b m = (6-0)/(2-4) = 6/-2 = -3 6 = (-3)(2) + b 6 = -6 + b b = 12 y = -3x + 12 y = (-3)(x - 4)

OpenStudy (anonymous):

Thank you!(: